Please help me with a simple solution. Help for Primary 5 maths.
14 years ago, when Mina celebrated her birthday with her family, the ratio of the age of Mina to Mina's sister was 7:9. This year, the ratio of the age of Mina to Mina's sister is 7:8. How old is Mina now?

Answer by Boreal(15235)

The ratio of their ages now is 7x/8x, so we need to find x
14 years ago ( 7x-14)/(8x-14) the ratio was 7/9
This is a proportion problem solved with cross multiplication.
(7x-14) *9=(8x-14) *7
Distribute
63x-126=56x -98
solve by first moving the x over, subtracting 56x from both sides
7x-126=-98
add 126 to both sides
7x=28
x=4
The ratio is 7x/8x, so Mina is 28. That is the answer asked for.
Her sister is now 32
14 years ago, Mina was 14 and her sister 18. That is a 7:9 ratio.